how does the time required for a first order reaction to go to 99% completion related to the half life of the reaction?


 

Grade:12th Pass

1 Answers

Aman Bansal
12 years ago

Dear Sajid,

The integrated rate law of first oder order reaction, which gives the change of concentration with time, is:
ln[A] = -k·t + ln[A]0
where [A]0 is the initial concentration.
Since ln(a) - ln(b) = ln(a/b) you can write
ln( [A]/[A]0 ) = -k·t

Next introduce the the fractional completion of the reaction
X = ( [A] - [A]0)/[A]0 = 1 - [A]/[A]0
So the integrated rate law in term of completion is:
ln(1 - X) = -k·t

From the given completion data you can calculate the rate constant:

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert